OpenStudy (anonymous):

I need help on this problem. How do I solve ln x + ln6x = 8 ? without a calculator ?

Directrix (directrix):

ln x + ln6x = 8 ln(x*(6x)) = 8 --> Sum of logs is log of product ln(6x^2) = 8 e^8 = 6 x^2 --> Definition of log e^8/6 = x^2 x^2 = (e^8)/6 x = (e^4) / Sqrt (6) --> Square root of both sides

OpenStudy (anonymous):

hmmm and that would be the final answer ?

Directrix (directrix):

x = (e^4) / Sqrt (6) is the exact answer. You can approximate for e and Sqrt(6) if that is what you mean.

OpenStudy (anonymous):

Alright thank you